OpenStudy (anonymous):

Happy sunday guys I need your help...taking a final test here is the quest... x/2-y/3=5/6 x/5-y/4=43/10 solve by substitution.....hugs

OpenStudy (anonymous):

I don't understand the substitution part. Do you mean simultaneous equations?

OpenStudy (anonymous):

yes i do hate it you isolate the y like I can solve one half but when i pkug in the second variable it does not match so 2x-3y=5

OpenStudy (anonymous):

-3y=5-2x y=-5/3+2/3x

OpenStudy (anonymous):

plug it intothe second equation and I get all kind of bizarre answers

OpenStudy (anonymous):

You can re-write the equations as 2x-3y=30 5x-4y=4300 This is simply just flipping the denominator so it multiplies with the numerator. Can you solve it now?

OpenStudy (anonymous):

I think so thanks for the idea:)

OpenStudy (anonymous):

Anymore problems just let me know :)

OpenStudy (anonymous):

:)